A restaurant uses five 80% energy efficient 30 W bulbs, a refrigerator marked “240 V, 5A” and a 1.5
kW electric cooker. These loads are connected to 220 V mains supply. The bulbs, refrigerator and
electric cooker are switch on for 12 hours, 24 hours and 5 hours respectively on daily average. The
restaurant opens six days a week – all loads are switched off throughout the 7th day.
(a) Determine the resistance of the refrigerator.
(b) (i) Calculate the electrical power in the refrigerator when plugged to the 220 V power outlet.
(ii) Calculate the input power of each light bulb.
(iii) At what rate is energy consumed by the restaurant?
(c) If the restaurant needs a single fuse to protect all its devices, recommend the rating of the fuse in
amperes.
(d) Calculate the daily energy consumption of the restaurant.
(e) If PHCN charges ₦32 per kWh, calculate the monthly (4 weeks) bill.
1
Expert's answer
2020-05-29T10:02:53-0400

(a) Determine the resistance of the refrigerator:


"R=\frac{V}{I}=\frac{240}{5}=48\space\Omega."

(b) (i) Calculate the electrical power in the refrigerator when plugged to the 220 V power outlet:


"P_r=\frac{V^2}{R}=\frac{220^2}{48}=1000\text{ W}."

(b) (ii) Calculate the input power of each light bulb:


"P_b=\frac{P_\text{rated}}{\eta}=\frac{30}{0.8}=37.5\text{ W}."

(b) (iii) At what rate is energy consumed by the restaurant?

The rate of energy consumption is energy per unit time, or power:

"P_\text{total}=5\cdot37.5+1000+1500=2690\text{ W}."

(c) If the restaurant needs a single fuse to protect all its devices, recommend the rating of the fuse in amperes is


"I=\frac{P_\text{total}}{V}=\frac{2690}{220}=12.2\text{ A}."


(d) Calculate the daily energy consumption of the restaurant.

The bulbs work 12 h, refrigerator works 24 h, electric cooker works 5 hours a day:


"E=5P_bt_b+P_rt_r+P_ct_c=\\
=5\cdot\frac{37.5}{1000}\cdot12+1\cdot24+1.5\cdot5=33.8\text{ kWh}."



(e) If PHCN charges ₦32 per kWh, calculate the monthly (4 weeks) bill.

The restaurant works six days a week, there are 4 weeks (24 days) in a month, the monthly bill is


"B_m=NEc=24\cdot33.8\cdot32=25958\text{ ₦/month}."